CC12:

What is the equation of a line that passes through the point (5, −4) and is parallel to the line whose equation is 2x + 5y = 10?

CC12:

@dude

CC12:

2x + 5y = 10 5y = -2x + 10 y = -2/5 x + 2 so slope = -2/5 passes through the point (5, −4) y = mx + b b = y - mx b = -4 - (-2/5)(5) b = -4 + 2 b = -2 equation y = -2/5 x - 2

CC12:

@AnimeGhoul8863 Can you help check my answer

AnimeGhoul8863:

urmmmmm e.e let meh think

AnimeGhoul8863:

i dun know T~T im sorry

CC12:

Its ok thank you